# Learning Goal-Conditioned Policies from Sub-Optimal Offline Data via Metric Learning

### v2

@@ A.3 Distance Monotonicity Measure
$\min_{\theta}$ $\operatorname{\mathbb{E}}_{D}\left[-\left\lVert\phi_{\theta}(s^{\prime\prime})-\phi_{\theta}(s)\right\rVert_{2}\right]$ (5)
[…]
Figure 7: The optimization in A.3 increases the distance between the representations. As long as $z_{2}$ is outside of the circumference, the triplet $s_{1},s_{2},s_{3}$ is distance monotonic.
This can be seen as the second term of the loss in Equation  1 with the first term as an explicit constraint. As defined before, the distance between any two points, $s_{i},s_{j}$, is the length of the geodesic on the graph defined as the offline dataset. This is equivalent to the sum of the intermediate steps within the geodesic path. Using the constraint in Equation A.3 we can use triangular inequality to bound the distance between points in the learned representation:
@@
A.3.1 Experiment on Distance Monotonicity Measure
@@
  • Concurrently we map these three points in $Z$ using the representation that’s being learned and compute the distances using a Euclidean norm, i.e., $d_{Z}(z_{1},z_{3})$ and $d_{Z}(z_{2},z_{3})$.
  • We then compute the distance monotonicity (DM) ratio as the number of triplets among these 1000 for which if $d_{G}(s_{1},s_{3})\leq d_{G}(s_{2},s_{3})$ then $d_{Z}(z_{1},z_{3})\leq d_{Z}(z_{2},z_{3})$ or if $d_{G}(s_{2},s_{3})\leq d_{G}(s_{1},s_{3})$ then $d_{Z}(z_{2},z_{3})\leq d_{Z}(z_{1},z_{3})$.
@@
Here we prove that, assuming a deterministic, sparse, goal-conditioned MDP, the greedy policy on a value function defined as in Equation 2 is equivalent to the greedy policy on the optimal value function.
To simplify the notation we will refer to $\phi(s)$ as $z$.
@@
$\pi^{\tilde{V}}_{g}(s)=\pi^{V^{*}}_{g}(s)$ (8)
$\implies$ $\operatorname*{arg\,max}_{a}\left[\tilde{V}(s^{\prime}=T(s,a))\right]=\operatorname*{arg\,max}_{a}\left[V^{*}(s^{\prime}=T(s,a))\right],$ (9)
here we make use of the fact that the MDP is deterministic and thus the tuple $(s,a)$ injectively corresponds to a unique next state $s^{\prime}$ through the transition function $T$.
We can rewrite the left-hand side using the definition of the value function of Equation 2 and the right-hand side using Bellman:
$\operatorname*{arg\,max}_{a}\left[\gamma^{d_{Z}(z^{\prime},z_{g})}r_{g}\right]=\operatorname*{arg\,max}_{a}\left[\max_{a^{\prime}}\left(r(s^{\prime},a^{\prime})+\gamma V^{*}(s^{\prime\prime})\right)\right],$ (10)
where again we denote $s^{\prime\prime}=T(s^{\prime},a^{\prime})$. Assuming a deterministic sparse goal-conditioned MDP, the maximum of the right-hand side in Equation 10 is equal to the discount factor raised to the power of the minimal number of actions needed to reach the goal, $d_{S}(s^{\prime},s_{g})$, times the reward at the goal, $r_{g}$:
$\max_{a^{\prime}}\left(r(s^{\prime},a^{\prime})+\gamma V^{*}(s^{\prime\prime})\right)=\gamma^{d_{S}(s^{\prime},s_{g})}r_{g}.$ (11)
We can now rewrite Equation 10 as:
$\operatorname*{arg\,max}_{a}\left[\gamma^{d_{Z}(z^{\prime},z_{g})}r_{g}\right]=\operatorname*{arg\,max}_{a}\left[\gamma^{d_{S}(s^{\prime},s_{g})}r_{g}\right],$ (12)
and change the argmax to an argmin by dropping the discount factor ($\gamma<1$) and the reward value:
$\operatorname*{arg\,min}_{a}\left[d_{Z}(z^{\prime},z_{g})\right]=\operatorname*{arg\,min}_{a}\left[d_{S}(s^{\prime},s_{g})\right].$ (13)
We can now make use of the distance monotonicity assumption of $\phi$. The action $a$ that minimizes the right-hand side of the above equation implies that for every other state $\hat{s}$ that can be reached from $s$ we have:
$d_{S}(s^{\prime},s_{g})\leq d_{S}(\hat{s},s_{g}),$ (14)
if $\phi$ is distance monotonic, this implies:
$d_{Z}(z^{\prime},z_{g})\leq d_{Z}(\hat{z},z_{g}),$ (15)
for their corresponding latent representation and thus the action that minimizes $d_{S}(s^{\prime},s_{g})$ is the same action that minimizes $d_{Z}(z^{\prime},z_{g})$ resulting in the equivalency:
$a_{\tilde{V}}=a_{V^{*}}\quad\forall s\in S,$ (16)
where we denoted $a_{\tilde{V}}$ as the action that minimizes $d_{Z}(z^{\prime},z_{g})$ and $a_{V^{*}}$ as the action that minimizes $d_{S}(s^{\prime},s_{g})$. This, in turn, implies that the two policies are equivalent thus concluding the proof. ∎
A.5 Incorporating Super-States in the Dataset
@@ A.6 Experimental Details
For each experiment, we provide the results for 5 runs with different seeds. Each model is trained for 100 epochs consisting of 500 batches of 256 data points each. Every model is trained using the Adam optimizer with a learning rate of $310^{-}$ 3. All experiments have been conducted using an NVIDIA RTX 3080 GPU accelerator.
@@ A.6.1 Model Hyperparameters
CQL: $\gamma=0.95$, we use a conservative weight of $5.0$.
BCQ: $\gamma=0.95$, action flexibility: 0.5, we sample $10$ actions per step and use 2 critics.
BEAR: $\gamma=0.95$, we use an adaptive $\alpha$ with an initial value of $0.001$ and a threshold of $0.05$, we use 2 critics modules and sample $10$ actions per step.
PLAS: $\gamma=0.95$, we use 2 critics modules.
IQL: $\gamma=0.95$ , we use 2 critics modules and an expectile value of $0.9$.
ContrastiveRL: $\gamma=0.95$, an offline regularization of $0.05$ and a fixed variance for the policy of $1$.
QRL: $\epsilon=0.25$, initial $\lambda=0.01$, offset softplus $500$ and $\beta=0.01$ and an offline regularization of $0.05$.
GoFAR: $\gamma=0.98$ and a discriminator gradient penalty of $0.01$. For the f-divergence we use the $\chi^{2}$-divergence.
@@ A.6.2 Evaluation Scenarios
Maze2D: The goal of the environment is to navigate an agent (actuated ball) to a target position inside a maze of varying complexity. The state space is 4-dimensional consisting of position and velocity in the plane, the action space is the torque applied to the ball in the two directions. The reward function is defined to be 1 when the agent reaches a position within 5 cm from the goal state and zero otherwise. Transitions take into consideration the momentum of the ball, frictions in the environment and collisions with the walls. For the low dataset we collect a dataset of 1000 trajectories of the agent performing uniform random actions. For the medium dataset we collect actions according to an Ornstein-Uhlenbeck process with parameters: $\theta=0.1$ and $\sigma=0.2$. For the high dataset we rely on the Minari dataset provided by [31].
Reach: The task consists of moving the end-effector of a simulated 7-DoF arm to a desired position in 3D. The state space consists of positions and velocities in 3D of the end effector and gripper of the robot and the goal refers to the desired position of the end-effector and zero velocity. Actions are translations of the end-effector. We first train a PPO agent online to solve the task and collect datasets at different stages of the training to define different qualities.
@@
For the results in Figure 6 we fix the maze to be 4 dimensional with 20 cells per dimension. We collect the datasets by first training a DQN agent online to solve the task and then collect 3 datasets using an $\epsilon$-greedy policy with $\epsilon$ respectively of values $0.9$, $0.5$ and $0.1$.
@@
Minigrid: We experiment with two variations of the minigrid environment. The first is the Empty environment where an agent translates freely within a grid-like environment. The state is described by the 2D position of the agent and the actions are the 4 possible translation directions. The reward is 1 once a randomly selected cell is reached and 0 otherwise. The DoorKey environment introduces bottlenecks in the MDP. A wall is introduced in the center of the grid separating it into 2 rooms with a closed door in the middle. In the first room, a key is placed in a random position. The goal is to pick up the key, open the door and reach a goal cell in the other room. The state space is defined as the position in the grid of the agent plus the position of the key plus a binary value describing whether the door is open or not. Actions are translations in the grid plus a pick-up action that has an effect only when the agent is adjacent to the key plus an open door action that has an effect only when the agent has the key and is adjacent to the door. As before, The datasets are collected by training a DQN agent online to convergence and collecting the datasets using $\epsilon$-greedy strategy with the three different values for $\epsilon$. For the high-dimensional observations case we use images rendered by the environment as the states. These are 80 by 80 pixels with 3 color channels.
@@ A.7 Sample-Efficiency
A main advantage of MetricRL stems from the nature of the loss function. Temporal difference methods (e.g., DQN) are known for their inefficiency when the time horizon grows considerably [56]. On the other hand, MetricRL mitigates this issue by using neural networks to learn a representation of a metric space.
To validate our hypothesis, we compare MetricRL against DQN [34] in the Hypermaze environment, considering a variable number of dimensions and cells. This allows us to control for both the dimensionality of the action space and the size of the state space of the underlying MDP.
@@ A.9 Additional Benchmark Results
Maze2D: We additionally provide the results for the u-maze and the medium maze described in [31], (Figures 11 and 12). Results confirm the findings described in the paper. MetricRL consistently outperforms the baselines in the case of low and medium datasets.
Figure 11: Average reward returns on Maze2D u-maze with different types of datasets. All results are averaged over 5 randomly-selected seeds. Higher is better.
Figure 12: Average reward returns on Maze2D medium maze with different types of datasets. All results are averaged over 5 randomly-selected seeds. Higher is better.
Minigrid: We provide results of the Empty environment with states and images as input, (Figures 13 and 14).
Figure 13: Average reward returns on Minigrid Empty with different types of datasets. All results are averaged over 5 randomly-selected seeds. Higher is better.
Figure 14: Average reward returns on Minigrid Empty with images with different types of datasets. All results are averaged over 5 randomly-selected seeds. Higher is better.
Below (Figure 15) are additional results on the increase in distance monotonicity measure paired with an increase in reward for different mazes and datasets of Maze2D.

### v3

@@ A.2 Distance Monotonicity Measure
$\min_{\theta}$ $\mathbb{E}_{D}\left[-\left\lVert\phi_{\theta}(s^{\prime\prime})-\phi_{\theta}(s)\right\rVert_{2}\right]$ (5)
$\text{subject to:}\left\lVert\phi_{\theta}(s^{\prime})-\phi_{\theta}(s)\right\rVert_{2}=1.$ (6)
Figure 7: The optimization in A.2 increases the distance between the representations. As long as $z_{2}$ is outside of the circumference, the triplet $s_{\{1,2,3\}}$ is distance monotonic.
This can be seen as the second term of the loss in Equation 1 with the first term as an explicit constraint. As defined before, the distance between any two points, $s_{i},s_{j}$, is the length of the geodesic on the graph defined as the offline dataset. This is equivalent to the sum of the intermediate steps within the geodesic path. Using the constraint in Equation A.2 we can use triangular inequality to bound the distance between points in the learned representation:
@@
A.2.1 Experiment on Distance Monotonicity Measure
@@
  • Concurrently we map these three points in $Z$ using the representation that’s being learned and compute the distances using a Euclidean norm, i.e., $d_{Z}(\phi(s_{1}),\phi(s_{3}))$ and $d_{Z}(\phi(s_{2}),\phi(s_{3}))$.
  • We then compute the distance monotonicity (DM) ratio as the number of triplets among these 1000 for which if $d_{G}(s_{1},s_{3})<d_{G}(s_{2},s_{3})$ then $d_{Z}(\phi(s_{1}),\phi(s_{3}))<d_{Z}(\phi(s_{2}),\phi(s_{3}))$ or if $d_{G}(s_{2},s_{3})<d_{G}(s_{1},s_{3})$ then $d_{Z}(\phi(s_{2}),\phi(s_{3}))<d_{Z}(\phi(s_{1}),\phi(s_{3}))$.
@@
Here we prove that, assuming a deterministic, sparse, goal-conditioned MDP, every optimal solution defined as the greedy policy based on the value function defined as in Equation 2 is contained in the greedy policy based on the optimal value function.
@@
$\pi_{g}^{\tilde{V}}(s)$ $=\operatorname*{arg\,max}_{a}\tilde{V}(s^{\prime}=T(s,a)),$ (8)
$\pi_{g}^{V^{*}}(s)$ $=\operatorname*{arg\,max}_{a}V^{*}(s^{\prime}=T(s,a)).$ (9)
Given that the MDP is deterministic, the transition function $T(s,a)$ injectively maps $(s,a)$ to a unique next state $s^{\prime}$.
From the definition of $\tilde{V}$ and the Bellman equation, we have:
$\tilde{V}(s^{\prime})$ $=\gamma^{d_{Z}(\phi(s^{\prime}),\phi(s_{g}))}r_{g},$ (10)
$V^{*}(s^{\prime})$ $=\max_{a^{\prime}}\left[r(s^{\prime},a^{\prime})+\gamma V^{*}(s^{\prime\prime})\right],$ (11)
where $s^{\prime\prime}=T(s^{\prime},a^{\prime})$. In a sparse, goal-conditioned MDP, the reward $r(s^{\prime},a^{\prime})$ is zero unless $s^{\prime}=s_{g}$, $V^{*}(s^{\prime})$ is equal to the discount factor raised to the power of the minimal number of actions needed to reach the goal and thus simplifies to:
$V^{*}(s^{\prime})=\gamma^{d_{S}(s^{\prime},s_{g})}r_{g}.$ (12)
Substituting the value functions into the greedy policy definitions, we get:
$\operatorname*{arg\,max}_{a}\gamma^{d_{Z}(\phi(s^{\prime}),\phi(s_{g}))}r_{g}\subseteq\operatorname*{arg\,max}_{a}\gamma^{d_{S}(s^{\prime},s_{g})}r_{g}.$ (13)
Since $\gamma<1$ and $r_{g}$ is constant, this reduces to:
$\operatorname*{arg\,min}_{a}d_{Z}(\phi(s^{\prime}),\phi(s_{g}))\subseteq\operatorname*{arg\,min}_{a}d_{S}(s^{\prime},s_{g}).$ (14)
[…]
$d_{Z}(\phi(\hat{s}),\phi(s_{g}))\leq d_{Z}(\phi(s^{\prime}),\phi(s_{g})),\quad\forall s^{\prime}\in T(s,\cdot),$ (15)
$d_{S}(\hat{s},s_{g})>d_{S}(\check{s},s_{g}),\quad\text{for some }\check{s}\in T(s,\cdot).$ (16)
By the distance monotonicity of $\phi$, $d_{S}(\hat{s},s_{g})>d_{S}(\check{s},s_{g})$ implies $d_{Z}(\phi(\hat{s}),\phi(s_{g}))>d_{Z}(\phi(\check{s}),\phi(s_{g}))$}, which contradicts the assumption that $\hat{a}\in\pi_{g}^{\tilde{V}}(s)$.
Thus, the actions minimizing $d_{Z}(\phi(s^{\prime}),\phi(s_{g}))$ are contained in the actions minimizing $d_{S}(s^{\prime},s_{g})$, implying:
$\pi_{g}^{\tilde{V}}(s)\subseteq\pi_{g}^{V^{*}}(s),\quad\forall s\in S.$ (17)
This completes the proof. ∎
A.4 Incorporating Super-States in the Dataset
@@
A.5 Pseudocode
We provide the pseudocode of our approach below.
Algorithm 1 MetricRL.
1:Initialize $\theta,\psi$
2:Offline dataset $B$, hyperparameters $\lambda$, $\eta$
3:repeat
4:  Sample batch $(s,a,s^{\prime},s_{g})_{\times B}\sim D$
5:  $s_{r}=\texttt{shuffle}(s)$ $\triangleright$ shuffle states in the batch
6:  $\mathcal{L}_{\theta}=(\lVert\phi_{\theta}(s)-\phi_{\theta}(s^{\prime})\rVert_{2}-1)^{2}-\lambda\log(\lVert\phi_{\theta}(s)-\phi_{\theta}(s_{r})\rVert_{2})$
7:  Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\theta}$
8:  $V=\lVert\phi_{\theta}(s)-\phi_{\theta}(s_{g})\rVert_{2}$
9:  $V^{\prime}=\lVert\phi_{\theta}(s^{\prime})-\phi_{\theta}(s_{g})\rVert_{2}$
10:  $\mathcal{L}_{\pi}=-(V^{\prime}-V)\log(\pi_{\psi}(a|s))$
11:  Update $\psi\leftarrow\psi-\eta\nabla_{\psi}\mathcal{L}_{\pi}$
12:until convergence
@@ A.6 Experimental Details
For each experiment, we provide the results for 5 runs with different seeds. Each model is trained for 100 epochs consisting of 500 batches of 256 data points each. Every model is trained using the Adam optimizer with a learning rate of $10^{-3}$. All experiments have been conducted using an NVIDIA RTX 3080 GPU accelerator.
@@ A.6.1 Models Description and Hyperparameters
CQL (Kumar et al., 2020), which introduces a conservative term in the estimation of the Q value. This term penalizes the highest values of the estimated Q function; hyperparameters: $\gamma=0.95$, we use a conservative weight of $5.0$.
BCQ (Fujimoto et al., 2019), which perturbs the policy learned with a VAE with a DDPG term; hyperparameters: $\gamma=0.95$, action flexibility: 0.5, we sample $10$ actions per step and use 2 critics.
BEAR (Kumar et al., 2019), that constrains the learned policy to the behavioral one estimated with BC; hyperparameters: $\gamma=0.95$, we use an adaptive $\alpha$ with an initial value of $0.001$ and a threshold of $0.05$, we use 2 critics modules and sample $10$ actions per step.
PLAS (Zhou et al., 2021), which trains the policy within the latent space of a conditional VAE trained on the Offline dataset; hyperparameters: $\gamma=0.95$, we use 2 critics modules.
IQL (Kostrikov et al., 2021), that avoids sampling out-of-distribution actions using a SARSA like critic update with quantile regression; hyperparameters: $\gamma=0.95$ , we use 2 critics modules and an expectile value of $0.9$.
ContrastiveRL (Eysenbach et al., 2022), which approximates the value function of the policy that collected the dataset using contrastive learning. To adapt it to offline RL, the objective is coupled with a behavioral cloning term. hyperparameters: $\gamma=0.95$, an offline regularization of $0.05$ and a fixed variance for the policy of $1$.
QRL (Wang et al., 2023), that estimates a quasimetric to approximate the value function using a contrastive learning formulation. The model is paired with a learned policy regularized with a behavioral cloning term to avoid out-of-distribution state-actions pairs in offline RL conditions. hyperparameters: $\epsilon=0.25$, initial $\lambda=0.01$, offset softplus $500$ and $\beta=0.01$ and an offline regularization of $0.05$.
GoFAR (Ma et al., 2022b), that estimates an offline goal-conditioned RL policy by recasting it as a state occupancy matching problem. hyperparameters: $\gamma=0.98$ and a discriminator gradient penalty of $0.01$. For the f-divergence we use the $\chi^{2}$-divergence.
HIQL (Park et al., 2023a), which reformulates IQL to an action-free hierarchical model. hyperparameters: $\gamma=0.99$ $\beta=1.0$, we use 2 critics modules and an expectile value of $0.7$.
@@ A.6.2 Evaluation Scenarios
Maze2D: The goal of the environment is to navigate an agent (actuated ball) to a target position inside a maze of varying complexity. The state space is 4-dimensional consisting of position and velocity in the plane, the action space is the torque applied to the ball in the two directions. The reward function is defined to be 1 when the agent reaches a position within 5 cm from the goal state and zero otherwise. Transitions take into consideration the momentum of the ball, frictions in the environment and collisions with the walls. For the low dataset we collect a dataset of 1000 trajectories of the agent performing uniform random actions, on average around 5% of successful trajectories. For the medium dataset we collect actions according to an Ornstein-Uhlenbeck process with parameters: $\theta=0.1$ and $\sigma=0.2$, on average 10% of successful trajectories. For the high dataset we rely on the Minari dataset provided by Younis et al. (2024), 99% successful trajectories.
Reach: The task consists of moving the end-effector of a simulated 7-DoF arm to a desired position in 3D. The state space consists of positions and velocities in 3D of the end effector and gripper of the robot and the goal refers to the desired position of the end-effector and zero velocity. Actions are translations of the end-effector. We first train a PPO agent online to solve the task and collect datasets at different stages of the training to define different qualities. Specifically, low has 20% successful trajectories, medium 60% and high 90%.
@@
For the results in Figure 6 we fix the maze to be 4 dimensional with 20 cells per dimension. We collect the datasets by first training a DQN agent online to solve the task and then collect 3 datasets using an $\epsilon$-greedy policy with $\epsilon$ respectively of values $0.9$ (success rate 2%), $0.5$ (success rate 60%) and $0.1$ (success rate 99%).
@@
Minigrid: We experiment with two variations of the minigrid environment. The first is the Empty environment where an agent translates freely within a grid-like environment. The state is described by the 2D position of the agent and the actions are the 4 possible translation directions. The reward is 1 once a randomly selected cell is reached and 0 otherwise. The DoorKey environment introduces bottlenecks in the MDP. A wall is introduced in the center of the grid separating it into 2 rooms with a closed door in the middle. In the first room, a key is placed in a random position. The goal is to pick up the key, open the door and reach a goal cell in the other room. The state space is defined as the position in the grid of the agent plus the position of the key plus a binary value describing whether the door is open or not. Actions are translations in the grid plus a pick-up action that has an effect only when the agent is adjacent to the key plus an open door action that has an effect only when the agent has the key and is adjacent to the door. As before, The datasets are collected by training a DQN agent online to convergence and collecting the datasets using $\epsilon$-greedy strategy with the same three different values for $\epsilon$ of before. The success rate is respectively of 70%, 100% and 100% for the Empty environment and 10%, 100% and 100% for the DoorKey environment. Notice how the main difference between the medium and high datasets here is not in the success rate but rather in the quality of the trajectories. For the high-dimensional observations case we use images rendered by the environment as the states. These are 80 by 80 pixels with 3 color channels.
@@ A.7 Sample-Efficiency
A main advantage of MetricRL stems from the nature of the loss function. Temporal difference methods (e.g., DQN) are known for their inefficiency when the time horizon grows considerably (Moore and Atkeson, 1993). On the other hand, MetricRL mitigates this issue by using neural networks to learn a representation of a metric space.
To validate our hypothesis, we compare MetricRL against DQN (Mnih et al., 2015) in the Hypermaze environment, considering a variable number of dimensions and cells. This allows us to control for both the dimensionality of the action space and the size of the state space of the underlying MDP.
@@ A.9 Additional Results
Value Function Estimation in Maze2D Large: Following the discussion of Section 3.2, we explore value function estimation in Maze2D Large for different types of datasets. The results of Figure 11 highlight that MetricRL is the only method able to correctly estimate the low value of the bottom left corner of the maze when the goal is on the other branch of the maze, regardless of the quality of the policy used to collect the offline dataset.
Figure 11: Estimated value function for different methods using a dataset collected from different policies in Maze2D Large. All values are normalized.
Maze2D: We additionally provide the results for the u-maze and the medium maze described in Fu et al. (2020), (Figures 12 and 13). Results confirm the findings described in the paper. MetricRL consistently outperforms the baselines in the case of low and medium datasets.
Figure 12: Average reward returns on Maze2D u-maze with different types of datasets. All results are averaged over 5 randomly selected seeds. Higher is better.
Figure 13: Average reward returns on Maze2D medium maze with different types of datasets. All results are averaged over 5 randomly selected seeds. Higher is better.
Minigrid: We provide results of the Empty environment with states and images as input, (Figures 14 and 15).
Figure 14: Average reward returns on Minigrid Empty with different types of datasets. All results are averaged over 5 randomly selected seeds. Higher is better.
Figure 15: Average reward returns on Minigrid Empty with images with different types of datasets. All results are averaged over 5 randomly selected seeds. Higher is better.
Below (Figure 16) are additional results on the increase in distance monotonicity measure paired with an increase in the reward for different mazes and datasets of Maze2D.